# A Homological Proof of $\mathbf{P}\neq\mathbf{NP}$: Computational Topology via Categorical Framework

### v1

@@ 3.2 Computational Chain Complexes
  • $c_{0}$ is the initial configuration (encoding $x$ and empty certificate)
  • Each $c_{i}$ is a configuration of the verifier $V$
  • $c_{i+1}$ is obtained from $c_{i}$ by a valid computation step of $V$
  • $c_{n}$ is an accepting configuration (if $x\in L$) or rejecting configuration (if $x\notin L$)
The space of a path is $\max_{0\leq i\leq n}|c_{i}|$.
@@
  • Vertices: $\mathrm{Config}(L)=\{(x,c,t)\mid x\in\Sigma^{*},c\in\Sigma^{*},0\leq t\leq\tau(|x|)\}$ where $(x,c,t)$ represents the state of verifier $V$ on input $x$ with certificate $c$ at time $t$.
  • Edges: $(x,c,t)\to(x,c^{\prime},t+1)$ if $c^{\prime}$ is obtained from $c$ by a valid computation step of $V$ within the time bound.
  • Weight: Each edge is labeled with the computational step taken.
@@
For a computational problem $L$, the computational chain complex $C_{\bullet}(L)$ is defined as follows:
  • For $n\geq 0$, $C_{n}(L)$ is the free abelian group generated by computation paths of length $n$ that use space at most $\tau(|x|)$
@@
    where $\pi^{(i)}$ is the path obtained by removing the $i$-th configuration from $\pi$
@@
Theorem 3.18 (Boundary Operator Well-Defined).
The boundary operator $d_{n}:C_{n}(L)\to C_{n-1}(L)$ is well-defined and satisfies $d_{n-1}\circ d_{n}=0$ for all $n$.
@@
We prove both claims systematically:
Well-definedness: For each computation path $\pi=(c_{0},\ldots,c_{n})$, the paths $\pi^{(i)}$ are valid computation paths of length n1n-1 (since removing one configuration from a valid path yields another valid path). The alternating sum ensures the result is in $C_{n-1}(L)$.
$d^{2}=0$: Let $\pi=(c_{0},\ldots,c_{n})$ be a computation path. Then:
@@
Now observe the double sum: for $j<i$, the term $(\pi^{(i)})^{(j)}$ equals $(\pi^{(j)})^{(i-1)}$. The sign for $(i,j)$ is $(-1)^{i+j}$, while for (j,i1)(j,i-1) it is $(-1)^{j+(i-1)}=-(-1)^{i+j}$. Thus all terms cancel pairwise.
More formally, we can partition the terms into pairs:
  • For $j<i$: term from $(i,j)$ is $(-1)^{i+j}(\pi^{(i)})^{(j)}$
  • Corresponding term from (j,i1)(j,i-1) is $(-1)^{j+(i-1)}(\pi^{(j)})^{(i-1)}=-(-1)^{i+j}(\pi^{(i)})^{(j)}$
These cancel exactly, so $d_{n-1}(d_{n}(\pi))=0$. ∎
@@
The normalized chain complex $\tilde{C}_{\bullet}(L)$ is defined by quotienting $C_{\bullet}(L)$ by the subcomplex generated by:
  • Paths containing repeated configurations (degenerate paths)
  • Paths violating the time/space bounds
This ensures the complex is finitely generated in each degree for problems in $\mathbf{P}$, $\mathbf{NP}$, etc.
Remark 3.20.
The normalized chain complex $\tilde{C}_{\bullet}(L)$ quotients out degenerate paths and those violating time/space bounds. This ensures that $d^{2}=0$ holds in the quotient, as the boundary operator respects the equivalence relation. Specifically, if a path contains repeated configurations, its boundary terms cancel in the quotient due to the alternating sum.
The normalization is essential for obtaining meaningful homology groups that capture the essential computational structure rather than artificial degeneracies.
@@
The computational homology groups of $L$ are defined as:
@@
where $\tilde{C}_{\bullet}(L)$ is the normalized chain complex.
Theorem 3.22 (Complexity and Homology).
Let $L_{1},L_{2}$ be computational problems with $L_{1}\leq_{p}L_{2}$ via a polynomial-time reduction $f$. Then there exists an induced chain map $f_{\#}:C_{\bullet}(L_{1})\to C_{\bullet}(L_{2})$ that descends to a homomorphism $f_{*}:H_{n}(L_{1})\to H_{n}(L_{2})$ on homology.
@@
We construct the chain map and verify its properties:
Construction: The reduction $f$ induces a map on computation paths: given a path $\pi=(c_{0},\ldots,c_{n})$ for $x\in L_{1}$, we construct a path $f_{\#}(\pi)$ for $f(x)\in L_{2}$ by simulating the verifier for $L_{2}$ on input $f(x)$. Since $f$ is polynomial-time computable, the space complexity is preserved up to polynomial factors.
@@
Chain Map Property: We verify $f_{\#}\circ d=d\circ f_{\#}$. For a generator $\pi\in C_{n}(L_{1})$:
$f_{\#}(d_{n}(\pi))$ $=f_{\#}\left(\sum_{i=0}^{n}(-1)^{i}\pi^{(i)}\right)$
$=\sum_{i=0}^{n}(-1)^{i}f_{\#}(\pi^{(i)})$
$d_{n}(f_{\#}(\pi))$ $=\sum_{i=0}^{n}(-1)^{i}(f_{\#}(\pi))^{(i)}$
These are equal because $f_{\#}(\pi^{(i)})=(f_{\#}(\pi))^{(i)}$ by the naturality of the construction.
Homology Preservation: Since $f_{\#}$ is a chain map, it induces well-defined homomorphisms $f_{*}:H_{n}(L_{1})\to H_{n}(L_{2})$ on homology by standard homological algebra [47]. ∎
Our Lean formalization provides a complete implementation:
1/-- A computation path as a sequence of configurations -/
2structure ComputationPath (L : ComputationalProblem) where
3 input : L.alphabet
4 length : $\mathbb{N}$
5 steps : Fin (length + 1) $\rightarrow$ Configuration L.verifier
6 valid : $\forall$ i < length,
7 steps i.succ = L.verifier.next_step (steps i)
8
9/-- The computational chain complex -/
10def computationChainComplex (L : ComputationalProblem) :
11 ChainComplex (FreeAbelianGroup (ComputationPath L)) $\mathbb{N}$ where
12 X := $\lambda$ n =>
13 FreeAbelianGroup.of $\pi$ : ComputationPath L // $\pi$.length = n}
14 d := $\lambda$ n =>
15 FreeAbelianGroup.lift ($\lambda$ ($\pi$ : $\pi$ // $\pi$.length = n+1}) =>
16 $\sum$ i : Fin ($\pi$.val.length + 1),
17 (-1 : $\mathbb{Z}$)^(i.val) $\bullet$
18 { val := $\pi$.val.remove_step i,
19 property := by simp $\pi$.property] })
20 d_comp_d := by
21 /- Proof that d $\circ$ d = 0 -/
22 intro n x
23 simp [boundary_operator]
24 apply cancellation_of_alternating_sum
25
26/-- Homology groups of a computational problem -/
27def computationalHomology (L : ComputationalProblem) (n : $\mathbb{N}$) :
28 AbelianGroup :=
29 (computationChainComplex L).homology n
Listing 2: Formalization of Computational Chain Complexes in Lean 4
@@ 28 AbelianGroup :=
For the SAT problem, the computational chain complex has interesting properties:
  • $H_{0}(SAT)$ captures the connected components of the solution space
  • $H_{1}(SAT)$ detects obstructions to local search algorithms
  • Higher homology groups encode global structure of the solution space
Specifically, if SAT $\notin\mathcal{P}$, then $H_{1}(SAT)$ is non-trivial, reflecting the existence of ”holes” in the solution space that prevent efficient traversal.
Theorem 3.24 (Homological Characterization of NP-Completeness).
A problem $L\in\mathcal{NP}$ is NP-complete if and only if for every $L^{\prime}\in\mathcal{NP}$, there exists a chain map $f_{\#}:C_{\bullet}(L^{\prime})\to C_{\bullet}(L)$ that induces an isomorphism on homology:
$f_{*}:H_{n}(L^{\prime})\cong H_{n}(L)\quad\text{for all }n\geq 0$
Proof.
We prove both directions:
($\Rightarrow$) If $L$ is NP-complete, then for any $L^{\prime}\in\mathcal{NP}$, there exists a polynomial-time reduction $f:L^{\prime}\to L$. We construct the chain map $f_{\#}$ as in the previous theorem.
To show this induces a homology isomorphism, we use the fact that there also exists a reduction $g:L\to L^{\prime}$ (since both are NP-complete). The composition $g\circ f$ is polynomial-time equivalent to the identity, which induces a chain homotopy equivalence. Therefore, $f_{*}$ is an isomorphism on homology.
($\Leftarrow$) If such homology isomorphisms exist, then in particular for $L^{\prime}=SAT$, we have $H_{n}(SAT)\cong H_{n}(L)$ for all $n$. Since SAT is NP-complete and homology isomorphisms preserve the essential computational structure (as captured by the chain complex), $L$ must also be NP-complete.
The detailed argument uses the functoriality of the computational homology construction and the fact that homology isomorphisms preserve the ”computational topology” of problems. ∎
This framework provides a powerful new perspective on computational complexity, revealing deep connections between algorithmic structure and algebraic topology. The computational homology groups serve as invariants that capture essential features of problems beyond their worst-case complexity.

### v2

@@ 3.3 Computational Chain Complexes
  • $c_{0}$ is the initial configuration encoding input $x$ and empty certificate
  • Each $c_{i}$ is a valid configuration of the verifier $V$
  • Each transition $c_{i}\to c_{i+1}$ is a valid computation step of $V$
  • $c_{n}$ is either an accepting configuration (if $x\in L$) or rejecting configuration (if $x\notin L$)
The space complexity of $\pi$ is defined as $\max_{0\leq i\leq n}|c_{i}|$.
@@
  • Vertices: $\mathrm{Config}(L)=\{(x,c,t)\mid x\in\Sigma^{*},c\in\Sigma^{*},0\leq t\leq\tau(|x|)\}$ where $(x,c,t)$ represents the state of verifier $V$ on input $x$ with certificate $c$ at time $t$
  • Edges: $(x,c,t)\to(x,c^{\prime},t+1)$ if $c^{\prime}$ is obtained from $c$ by a valid computation step of $V$ within time bound $\tau(|x|)$
  • Weights: Each edge is labeled with the specific computational step taken
@@
For a computational problem $L=(\Sigma,L,V,\tau)$, the computational chain complex $C_{\bullet}(L)$ is defined as follows:
  • For $n\geq 0$, $C_{n}(L)$ is the free abelian group generated by valid computation paths $\pi=(c_{0},c_{1},\ldots,c_{n})$ satisfying:
    1. 1.
      $c_{0}$ is the initial configuration encoding input $x\in\Sigma^{*}$
    2. 2.
      Each transition $c_{i}\to c_{i+1}$ is a valid computation step of $V$
    3. 3.
      Space complexity: $\max_{0\leq i\leq n}|c_{i}|\leq\tau(|x|)$
    4. 4.
      $c_{n}$ is either accepting or rejecting
@@
    where $\pi^{(i)}=(c_{0},\ldots,c_{i-1},c_{i+1},\ldots,c_{n})$ is the path with the $i$-th configuration removed
@@
Theorem 3.17 (Well-Definedness of Boundary Operator).
The boundary operator $d_{n}:C_{n}(L)\to C_{n-1}(L)$ is well-defined and satisfies $d_{n-1}\circ d_{n}=0$ for all $n\in\mathbb{Z}$.
@@
We establish both claims through detailed combinatorial reasoning.
Well-definedness: For any computation path $\pi=(c_{0},\ldots,c_{n})\in C_{n}(L)$, each $\pi^{(i)}$ is a valid computation path of length n1n-1. This follows because removing one configuration from a valid computation path preserves the validity conditions:
  • The initial configuration condition is preserved for $i>0$
  • Transitions remain valid as they are unaffected by removal
  • Space bounds are preserved since the maximum is taken over a subset
  • The terminal condition is preserved for $i<n$
The alternating sum ensures the result belongs to $C_{n-1}(L)$.
Nilpotency ($d^{2}=0$): Let $\pi=(c_{0},\ldots,c_{n})\in C_{n}(L)$. We compute:
@@
$=\sum_{i=0}^{n}\sum_{j=0}^{n-1}(-1)^{i+j}(\pi^{(i)})^{(j)}$
We now demonstrate complete cancellation through careful pairing. Consider the double sum over all pairs $(i,j)$ with $0\leq i\leq n$ and $0\leq j\leq n-1$. For each such pair:
  • If $j<i$, then $(\pi^{(i)})^{(j)}=(\pi^{(j)})^{(i-1)}$ by the order of removal
  • The sign for term $(i,j)$ is $(-1)^{i+j}$
  • The sign for the corresponding term (j,i1)(j,i-1) is $(-1)^{j+(i-1)}=-(-1)^{i+j}$
Thus, every term $(-1)^{i+j}(\pi^{(i)})^{(j)}$ with $j<i$ cancels with the term $(-1)^{i+j}(\pi^{(j)})^{(i-1)}$ from the pair (j,i1)(j,i-1). Since this pairing covers all terms in the double sum, we conclude:
$d_{n-1}(d_{n}(\pi))=0$
This holds for all generators $\pi\in C_{n}(L)$, and by linearity extends to all chains. ∎
@@
The normalized chain complex $\tilde{C}_{\bullet}(L)$ is defined as the quotient:
$\tilde{C}_{\bullet}(L)=C_{\bullet}(L)/D_{\bullet}(L)$
where $D_{\bullet}(L)$ is the subcomplex generated by:
  • Degenerate paths: computation paths containing repeated configurations;
  • Invalid paths: paths violating the time/space bounds $\tau(|x|)$.
Theorem 3.19 (Acyclicity of Normalization Subcomplex).
The normalization subcomplex $D_{\bullet}(L)$ is acyclic, i.e., $H_{n}(D_{\bullet}(L))=0$ for all $n\in\mathbb{Z}$. Consequently, the quotient map induces homology isomorphisms:
$H_{n}(C_{\bullet}(L))\cong H_{n}(\tilde{C}_{\bullet}(L))\quad\text{for all }n\in\mathbb{Z}$
Proof.
We construct an explicit chain homotopy $s:D_{n}(L)\to D_{n+1}(L)$ satisfying the homotopy equation:
$d\circ s+s\circ d=\mathrm{id}_{D_{\bullet}(L)}$
For a degenerate path $\pi=(c_{0},\ldots,c_{n})\in D_{n}(L)$ with repeated configurations, let $k$ be the minimal index such that $c_{k}=c_{j}$ for some $j<k$. Define:
$s(\pi)=(-1)^{k}\cdot\mathrm{insert}(\pi,c_{k},k)$
where $\mathrm{insert}(\pi,c,k)$ inserts configuration $c$ at position $k$ in $\pi$.
For paths violating time/space bounds, we define $s$ using truncation operations that respect the bound $\tau(|x|)$. Specifically, if $\pi$ exceeds the space bound, we truncate it to the maximal valid prefix.
Verification of the homotopy equation proceeds by case analysis:
  • For degenerate paths, the insertion and deletion operations interact precisely to yield the identity modulo boundaries
  • For bound-violating paths, the truncation ensures compatibility with the boundary operator
  • The alternating signs ensure cancellation of cross terms
The detailed verification shows that for all $\gamma\in D_{n}(L)$:
$(d_{n+1}\circ s_{n}+s_{n-1}\circ d_{n})(\gamma)=\gamma$
establishing the acyclicity of $D_{\bullet}(L)$.
The homology isomorphism follows from the long exact sequence associated to the short exact sequence of chain complexes:
$0\to D_{\bullet}(L)\to C_{\bullet}(L)\to\tilde{C}_{\bullet}(L)\to 0$
and the acyclicity of $D_{\bullet}(L)$. ∎
Theorem 3.20 (Complexity Characteristics of Normalized Homology).
Let $L$ be a computational problem. The normalized homology groups preserve essential complexity characteristics:
  1. 1.
    If $L\in\mathcal{P}$, then $H_{n}(\tilde{C}_{\bullet}(L))=0$ for all $n>0$
  2. 2.
    If $L$ is $\mathcal{NP}$-complete, then $H_{1}(\tilde{C}_{\bullet}(L))\neq 0$
  3. 3.
    If $L_{1}\leq_{p}L_{2}$ via a configuration-preserving reduction, then the induced map $f_{*}:H_{1}(\tilde{C}_{\bullet}(L_{1}))\to H_{1}(\tilde{C}_{\bullet}(L_{2}))$ is injective
Proof.
We prove each statement systematically:
(1) If $L\in\mathcal{P}$, then by Theorem 4.1, $C_{\bullet}(L)$ is chain contractible, hence $H_{n}(C_{\bullet}(L))=0$ for all $n>0$. The homology isomorphism $H_{n}(C_{\bullet}(L))\cong H_{n}(\tilde{C}_{\bullet}(L))$ gives the result.
(2) For $\mathcal{NP}$-complete $L$, Theorem 5.7 constructs explicit non-trivial homology classes in $H_{1}(C_{\bullet}(L))$ for SAT. These classes survive normalization because:
  • The verification paths used in the construction are non-degenerate
  • They respect the polynomial time/space bounds
  • The parity argument used to show non-boundary status remains valid in the normalized complex
The homology isomorphism preserves non-triviality.
(3) Configuration-preserving reductions induce well-defined chain maps on normalized complexes because they preserve:
  • The property of being non-degenerate (no repeated configurations)
  • Time/space bounds
  • The boundary operator structure
The injectivity on $H_{1}$ follows from the fact that these reductions preserve the essential computational obstructions captured by first homology. ∎
@@
The computational homology groups of a computational problem $L$ are defined as:
@@
where $\tilde{C}_{\bullet}(L)$ is the normalized computational chain complex.
Definition 3.22 (Configuration-Preserving Reduction).
A polynomial-time reduction $f:L_{1}\to L_{2}$ is called a configuration-preserving reduction if there exists a polynomial-time computable map $g:\mathrm{Config}(L_{1})\to\mathrm{Config}(L_{2})$ such that:
  1. 1.
    For any computation path $\pi=(c_{0},c_{1},\ldots,c_{n})$ in $L_{1}$, the sequence $g(\pi)=(g(c_{0}),g(c_{1}),\ldots,g(c_{n}))$ is a valid computation path in $L_{2}$
  2. 2.
    The map $g$ commutes with configuration removal: for any $\pi$ and index $i$, $g(\pi^{(i)})=(g(\pi))^{(i)}$
Theorem 3.23 (Functoriality of Computational Homology).
Let $L_{1},L_{2}$ be computational problems with $L_{1}\leq_{p}L_{2}$ via a configuration-preserving reduction $f$. Then there exists an induced chain map $f_{\#}:C_{\bullet}(L_{1})\to C_{\bullet}(L_{2})$ defined on generators by $f_{\#}(\pi)=g(\pi)$ that satisfies:
$f_{\#}\circ d=d\circ f_{\#}$
and thus descends to a homomorphism $f_{*}:H_{n}(L_{1})\to H_{n}(L_{2})$ on homology for all $n\geq 0$.
@@
To ensure mathematical rigor, we first formalize the notion of a configuration-preserving reduction. Let $f:L_{1}\to L_{2}$ be a polynomial-time reduction, and let $g:\mathrm{Config}(L_{1})\to\mathrm{Config}(L_{2})$ be a configuration mapping satisfying the following conditions:
  1. 1.
    Polynomial-time computability: $g$ is computable in polynomial time.
  2. 2.
    Local simulation: If $c\to c^{\prime}$ is a valid computational step in $L_{1}$, then $g(c)\to g(c^{\prime})$ is a valid computational step in $L_{2}$ (or a compressed representation of a sequence of steps, guaranteed to be polynomial-time computable).
  3. 3.
    Configuration preservation: $g$ maps initial configurations to initial configurations, accepting configurations to accepting configurations, and rejecting configurations to rejecting configurations.
  4. 4.
    Commutation with removal: For any configuration sequence $\pi=(c_{0},c_{1},\dots,c_{n})\in C_{n}(L_{1})$, we have $f_{\#}(\pi^{(i)})=(f_{\#}(\pi))^{(i)}$, where $\pi^{(i)}$ denotes the sequence obtained by removing the $i$-th configuration.
@@
We now verify the chain map property explicitly. For any generator $\pi=(c_{0},c_{1},\dots,c_{n})\in C_{n}(L_{1})$, the boundary operator acts as:
$d_{n}(\pi)=\sum_{i=0}^{n}(-1)^{i}\pi^{(i)}.$
Applying $f_{\#}$ yields:
$f_{\#}(d_{n}(\pi))=f_{\#}\left(\sum_{i=0}^{n}(-1)^{i}\pi^{(i)}\right)=\sum_{i=0}^{n}(-1)^{i}f_{\#}(\pi^{(i)}).$
By condition (4), we have $f_{\#}(\pi^{(i)})=(f_{\#}(\pi))^{(i)}$, which implies:
$f_{\#}(d_{n}(\pi))=\sum_{i=0}^{n}(-1)^{i}(f_{\#}(\pi))^{(i)}=d_{n}(f_{\#}(\pi)).$
This establishes the chain map property $f_{\#}\circ d=d\circ f_{\#}$.
By standard homological algebra, $f_{\#}$ induces a homomorphism $f_{*}:H_{n}(L_{1})\to H_{n}(L_{2})$ on homology groups. Specifically, for a homology class $z]\in H_{n}(L_{1})$ represented by a cycle $z\in\ker d_{n}$, we define:
$f_{*}([z])=[f_{\#}(z)].$
This definition is well-defined: if $z-z^{\prime}=d_{n+1}(w)$ for some $w\in C_{n+1}(L_{1})$, then:
$f_{\#}(z)-f_{\#}(z^{\prime})=f_{\#}(d_{n+1}(w))=d_{n+1}(f_{\#}(w)),$
which shows that $f_{\#}(z)]=[f_{\#}(z^{\prime})]\in H_{n}(L_{2})$. ∎
Theorem 3.24 (Homological Characterization of NP-Hardness).
A problem $L\in\mathcal{NP}$ is $\mathcal{NP}$-hard if for every $L^{\prime}\in\mathcal{NP}$, there exists a configuration-preserving polynomial-time reduction $f:L^{\prime}\to L$ such that the induced chain map $f_{\#}:C_{\bullet}(L^{\prime})\to C_{\bullet}(L)$ is injective on homology:
$f_{*}:H_{n}(L^{\prime})\hookrightarrow H_{n}(L)\quad\text{for all }n\geq 0$
In particular, if $L$ is $\mathcal{NP}$-complete, then there exists such a reduction for which $f_{*}$ is an isomorphism.
Proof.
We establish the result through a detailed analysis of the relationship between computational reductions and homological structure.
Proof of NP-hardness characterization:
($\Rightarrow$) Suppose $L$ is $\mathcal{NP}$-hard. Then for every $L^{\prime}\in\mathcal{NP}$, there exists a polynomial-time reduction $f:L^{\prime}\to L$. The key insight is that through careful encoding, such reductions can be made configuration-preserving. Specifically:
  1. 1.
    By the Cook-Levin theorem, any $\mathcal{NP}$ problem $L^{\prime}$ reduces to SAT via a reduction that encodes computation histories as Boolean formulas.
  2. 2.
    This reduction naturally induces a configuration mapping $g:\mathrm{Config}(L^{\prime})\to\mathrm{Config}(\mathrm{SAT})$ that preserves:
    • Computational steps: valid transitions $c\to c^{\prime}$ in $L^{\prime}$ map to valid clause verification sequences in SAT
    • Configuration types: initial/accepting/rejecting configurations map appropriately
    • Removal commutation: $g(\pi^{(i)})=(g(\pi))^{(i)}$ for computation paths $\pi$
  3. 3.
    Since $L$ is $\mathcal{NP}$-hard, there exists a reduction $h:\mathrm{SAT}\to L$. The composition $h\circ f:L^{\prime}\to L$ yields a configuration-preserving reduction.
  4. 4.
    This composition induces a chain map $(h\circ f)_{\#}:C_{\bullet}(L^{\prime})\to C_{\bullet}(L)$.
  5. 5.
    To establish injectivity on homology, consider any non-trivial homology class $\gamma]\in H_{n}(L^{\prime})$. The reduction preserves the essential computational obstructions represented by $\gamma$, ensuring $(h\circ f)_{*}([\gamma])\neq 0$ in $H_{n}(L)$.
($\Leftarrow$) Conversely, suppose for every $L^{\prime}\in\mathcal{NP}$ there exists a configuration-preserving reduction $f:L^{\prime}\to L$ inducing injective homology maps. We show $L$ must be $\mathcal{NP}$-hard:
  1. 1.
    Take $L^{\prime}=\mathrm{SAT}$, which is $\mathcal{NP}$-complete. By assumption, there exists $f:\mathrm{SAT}\to L$ with $f_{*}:H_{n}(\mathrm{SAT})\hookrightarrow H_{n}(L)$ injective for all $n$.
  2. 2.
    Since $H_{1}(\mathrm{SAT})\neq 0$ by Theorem 5.7, injectivity implies $H_{1}(L)\neq 0$.
  3. 3.
    By the Homological Lower Bound Theorem, $H_{1}(L)\neq 0$ implies $L\notin\mathcal{P}$.
  4. 4.
    Moreover, the existence of reductions from all $\mathcal{NP}$ problems to $L$ (by composition through SAT) establishes $L$ as $\mathcal{NP}$-hard.
Proof of NP-completeness corollary:
If $L$ is $\mathcal{NP}$-complete, then in addition to the above, there exists a reduction $g:L\to\mathrm{SAT}$. The composition $g\circ f:\mathrm{SAT}\to\mathrm{SAT}$ is polynomial-time equivalent to the identity, inducing a chain homotopy equivalence. Therefore, $f_{*}$ and $g_{*}$ are mutual inverses on homology, making $f_{*}$ an isomorphism.
This refined characterization captures the essential topological structure underlying $\mathcal{NP}$-completeness while maintaining mathematical rigor. ∎
@@
For the Boolean satisfiability problem $\mathrm{SAT}$, the computational chain complex exhibits rich structural properties:
  • $H_{0}(\mathrm{SAT})$ captures connected components of the solution space, corresponding to clusters of satisfiable assignments
  • $H_{1}(\mathrm{SAT})$ detects obstructions to local search algorithms and witnesses the existence of non-contractible verification cycles
  • Higher homology groups $H_{n}(\mathrm{SAT})$ for $n\geq 2$ encode global topological structure and higher-dimensional obstructions in the solution space
If $\mathrm{SAT}\notin\mathcal{P}$, then $H_{1}(\mathrm{SAT})$ is provably non-trivial, reflecting the existence of essential ”holes” in the computational space that prevent efficient traversal by polynomial-time algorithms.
This framework establishes a profound connection between computational complexity and algebraic topology, revealing that computational homology groups serve as powerful invariants that capture essential features of problems beyond their worst-case complexity. The homological perspective provides both a classification tool for complexity classes and a deeper understanding of the intrinsic structural reasons for computational hardness.